OpenStudy (mindblast3r):

Samuel has a collection of toy cars. His favorites are the 27 red ones, which make up 60% of his collection. How many toy cars does Samuel have?

OpenStudy (mindblast3r):

I already know the answer and how to get it in cross multiplication form, but how do I write this question in algebraic form?

OpenStudy (mindblast3r):

the answer is 45.

OpenStudy (math&ing001):

Yep ! What do you mean by algebraic form ?

OpenStudy (mindblast3r):

\[\frac{ 27 }{ 60 } \frac{ x }{ 100 }\]

OpenStudy (mindblast3r):

X = 27 x 100 ÷ 60 X = 45

OpenStudy (mindblast3r):

how would i write the question into an algebraic form now? That's all I want to know if you know.

OpenStudy (mindblast3r):

oh i didn't read what you wrote. now i did.

OpenStudy (mindblast3r):

something like 27x = 60

OpenStudy (mindblast3r):

or something like that.

OpenStudy (mindblast3r):

not even sure how to write it.

OpenStudy (math&ing001):

I think you already wrote it somewhere above. It would look like (60/100)*x - 27 =0

OpenStudy (math&ing001):

It's not an emoticon at the end btw that's 'equal zero'.

OpenStudy (mindblast3r):

hm

OpenStudy (mindblast3r):

i think you're right.

OpenStudy (mindblast3r):

what if i did 60% ÷ 3 = 20% x 5 = 100% so 27 ÷ 3 x 5 = 45

OpenStudy (mindblast3r):

do you get what i did?

OpenStudy (math&ing001):

Nope.

OpenStudy (mindblast3r):

i turned the 27 (60%) into 20% then i multiplied it by 5 to get 100%, because 20 x 5 = 100%

OpenStudy (math&ing001):

A lot of numbers you introduced do not exist on the initial problem.

OpenStudy (math&ing001):

All i did was write the problem in another way 60% * x = 27 so 60% *x - 27 = 0 I don't get why you would do what you did.

OpenStudy (mindblast3r):

how did you do that? Like how did you know to write the problem like that?

OpenStudy (mindblast3r):

that is what I initially wanted to know lol.

OpenStudy (mindblast3r):

pls respond

OpenStudy (mindblast3r):

how did you know to write the problem in that way?

OpenStudy (math&ing001):

27 cars represent 60% of his collection, and x represents 100% of his collection. So 60% * x =27 that's how you get it.

OpenStudy (mindblast3r):

.60x = 27 .60 ÷ .60 = 1 27 ÷ .60 = 45 X = 45

OpenStudy (mindblast3r):

ahhh I get it now.

OpenStudy (math&ing001):

Writing in algebraic form means finding a function 'f' so that f(x)=0. That's why I wrote it that way.

OpenStudy (mindblast3r):

? function? lol your speaking to me in chinese

OpenStudy (math&ing001):

You haven't study functions yet ?

OpenStudy (mindblast3r):

no ><

OpenStudy (math&ing001):

studied*

OpenStudy (math&ing001):

Oh ok :P

OpenStudy (mindblast3r):

didn't learn that yet :P

OpenStudy (mindblast3r):

Is it hard?

OpenStudy (mindblast3r):

Right now, I'm just learning basic algebra.

OpenStudy (math&ing001):

Not really, no.

OpenStudy (mindblast3r):

Hmm..

OpenStudy (mindblast3r):

how far away am I from learning functions?

OpenStudy (mindblast3r):

wait, i don't even need to worry about that stuff right now lol.

OpenStudy (math&ing001):

Yeah :p and good luck n your studies !

OpenStudy (mindblast3r):

wait, so the way you wrote it the problem what is it called? like for example the first way I did it was called the cross multiplication method.

OpenStudy (mindblast3r):

wait, what do you mean good luck? you think i need luck? lol

OpenStudy (math&ing001):

I'm sure you don't :3

OpenStudy (mindblast3r):

what makes you think that?

OpenStudy (math&ing001):

Well you got the problem right the first time. Anyways I'm out, have a good day !

OpenStudy (mindblast3r):

ok thanks a lot bye :)

OpenStudy (math&ing001):

Welcome !